Methane molecules in 112cm3 of methane are standard temperature and pressure


1
Expert's answer
2021-01-29T08:24:26-0500

Volume = 112 cm3 = 112/1000 = 0.112 L

Standard Temperature and pressure = 273 K and 1 atm

Using Ideal gas equation,

PV = nRT

Where, P = pressure

V = Volume

n = moles

R = Gas constant = 0.0821atm-L/mol-K

T = Temperature

n = PV/RT

= 1.0×0.112/0.0821×273

= 0.0050 mol

Number of molecules = mole(n) × NA

NisAvogadro Number = 6.022×1023

Number of Methane molecules = 0.0050×6.022×1023

= 3.00×1021 molecules

[So, Methane molecules = 3.00×1021 molecules]